OpenStudy (anonymous):

I need to solve for X: 3(X-2)=-6 I would like to see how to do it, rather than just the answer! Thanks!

OpenStudy (anonymous):

3(x-2) = -6 Start by multiplying each of the terms in the parenthesis by 3. So.. 3*x = 3x 3 * -2 = -6 Now we have 3x - 6 = -6 To get x alone we have to add six to the left side... and because we do that we have to do the same to the right side. So on the left side we're left with 3x and on the right side we have 3x = 0 Now to just get x alone we have to divide both sides by the number sitting in front of the x which is 3. So... 3x/3 = 0/3 We end up with x=0 Hope that was clear enough!

OpenStudy (abb0t):

Showing you the steps to do it is the same as giving you the answer. Lol

OpenStudy (anonymous):

@abb0t Eh. Chances are they have more similar questions and showing them how to do one step by step will (hopefully) teach them how to do the rest on their own.

OpenStudy (anonymous):

@abb0t actually it is not the same. One is giving a fish, the other is teaching to fish so he can become self sufficient.

OpenStudy (anonymous):

Thanks IAmAUser42! That really helped!

OpenStudy (anonymous):

@Hello_Jello2468 Awesome! You're welcome